### v1

@@ Abstract
Identifying procedural errors online from egocentric videos is a critical yet challenging task across various domains, including manufacturing, healthcare, and skill-based training. The nature of such mistakes is inherently open-set, as unforeseen or novel errors may occur, necessitating robust detection systems that do not rely on prior examples of failure. Currently, however, no technique effectively detects open-set procedural mistakes online.
We propose a dual-branch architecture to address this problem in an online fashion: one branch continuously performs step recognition from the input egocentric video, while the other anticipates future steps based on the recognition module’s output. Mistakes are detected as mismatches between the currently recognized action and the action predicted by the anticipation module. The recognition branch takes input frames, predicts the current action, and aggregates frame-level results into action tokens. The anticipation branch, specifically, leverages the solid pattern-matching capabilities of Large Language Models (LLMs) to predict action tokens based on previously predicted ones.
Given the online nature of the task, we also thoroughly benchmark the difficulties associated with per-frame evaluations, particularly the need for accurate and timely predictions in dynamic online scenarios.
Extensive experiments on two procedural datasets demonstrate the challenges and opportunities of leveraging a dual-branch architecture for mistake detection, showcasing the effectiveness of our proposed approach. In a thorough evaluation—including recognition and anticipation variants and state-of-the-art models—our method reveals its robustness and effectiveness in online applications.
@@ 1 Introduction
Detecting procedural errors in videos has recently gained significant attention due to its potential to provide substantial benefits across various domains. Accurate mistake detection in videos of individuals performing tasks—such as recipe execution, object assembly, or managing complex workflows—holds promise for enhancing training and learning experiences by providing real-time feedback (see Figure 1). This immediate feedback allows for timely corrections, fostering faster skill development, skill acquisition, and a safer learning environment in high-stakes areas like surgery and aviation.
@@
Figure 1: Procedural Mistake Detection involves identifying errors within a procedural video, where each procedure consists of steps that can be correct or incorrect. The aim is to develop a method capable of analyzing a video and determining whether each frame contains a mistake. This concept is particularly useful when applied to egocentric videos, as they allow for direct feedback to the person performing the task. In the diagram, the leftmost column shows the steps that define a procedure, while the right side presents a sequence of frames, which can either indicate a mistake (red) or a correct action (green). The black arrow illustrates the flow of time, and the algorithm, depicted by a gear icon, takes the video sequence from time 00 to $t-1$ as input, classifying the current frame $t$ as either correct (green) or a mistake (red).
@@
Advanced mistake detection models will become essential for ensuring precision, safety, and efficiency in procedural applications. This demand has driven the creation of new datasets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and methodologies designed to advance procedure learning [13, 4, 14, 15, 1, 6, 16] and error detection models [3, 5, 8]. Currently, existing mistake detection approaches vary widely. Some methods emphasize action detection, identifying specific errors like missing steps or out-of-sequence actions [17], or incorrect oredering [16]. While other models bypass action detection altogether and monitor changes to the assembled object to ensure procedural accuracy [18, 19].
@@
An ideal Procedural Mistake Detection (PMD) model should demonstrate two critical capabilities: robustness to diverse mistake types and the ability to provide online, timely feedback. First, a mistake detection system should identify any deviation from the correct procedure, regardless of the error’s nature. Since the range of correct methods is finite while possible mistakes are virtually unlimited, we use the One-Class Classification (OCC) framework. OCC trains on correct procedure videos only and tests on a mix of correct and incorrect cases, similar to anomaly detection [20, 21]. Second, the model should provide online feedback quickly enough to allow users to correct mistakes promptly, minimizing the reinforcement of incorrect actions.  [^1] We distinguish an online setup, which allows slight processing delays while enabling effective feedback, from a real-time setup, which requires immediate responses. Additionally, an egocentric perspective enhances the model’s applicability by mirroring natural human perception, thereby increasing feasibility for real-world deployment in task execution scenarios.
@@
We propose a dual-branch approach combining step recognition and anticipation to detect mismatches in predicted and observed actions. The first branch, a step recognition module, processes the input video to identify the current action being performed. The second branch, a step anticipation module, predicts the next step based on previously recognized actions, utilizing a language model’s reasoning capabilities. Mistakes are flagged when there is a divergence between recognized and anticipated actions, highlighting inconsistencies between actual and expected steps.
This paper extends our previous conference work, PREGO [22], by refining key components to improve procedural mistake detection. While PREGO’s evaluation of Large Language Models (LLMs) focused on a single model and fixed approach, our work conducts a comprehensive evaluation of LLMs as step anticipators, achieving state-of-the-art results on Assembly101-O and Epic-tent-O [22]. Specifically, we examine a broader range of LLMs, fine-tune them using Low-Rank Adaptation (LLoRA), and experiment with various prompting strategies, including Zero-Shot, Few-Shot, and Automatic Chain of Thought (Auto-CoT) [23], with Auto-CoT significantly enhancing the anticipation branch’s predictive abilities.
Further, we address noise in the recognizer’s frame-by-frame predictions, which can disrupt action sequence continuity. To mitigate this, we extend beyond fixed-window aggregation to explore alternative strategies that reduce noise and improve prediction accuracy. We also introduce frame-level metrics for more precise evaluation, aligning the recognition module’s output with the anticipation module’s input. Given the online nature of this task, we emphasize accurate, timely per-frame evaluations in dynamic scenarios to ensure effective mistake detection and reduce response delays.
@@
Our contributions include (1) exploring and benchmarking multiple LLMs to advance PREGO’s foundational approach, (2) experimenting with various prompting techniques, including Auto-CoT, which enhances overall performance, (3) evaluating frame-level metrics for more precise system assessment compared to the original fixed window approach, and (4) investigating alternative aggregation techniques to improve mistake detection robustness and continuity. Extensive experimentation demonstrates the challenges and opportunities in dual-branch architectures and LLM-based step anticipation for open-set mistake detection. Our results underscore the potential of integrating step recognition and anticipation within a unified framework for online procedural error detection, offering a more flexible and effective solution for real-world applications.
@@ 2.1 Procedural Mistake Detection
Procedural learning has made notable progress with the development of diverse datasets [6, 10, 9, 11, 12], offering valuable insights into both structured [7, 3, 24] and unstructured [1, 25] tasks. These datasets span a wide range of applications, from industrial assembly [24, 2, 8, 12] to everyday cooking activities [25, 26, 27]. Despite the growing interest in this field, the lack of a standardized method for mistake detection has resulted in limited evaluation and fragmented literature.
@@ 2.2 Datasets of Procedural Egocentric Videos
Assembly101 [2] is a large-scale video dataset that provides frame-level mistake annotations. It features videos of actors assembling toys, with synchronized Ego-Exo views and annotated hand positions. IndustReal [8] focuses on a single toy, which leads to a single procedure being learned. Epic-tent [1] is a dataset that covers a different domain of unscripted actions that capture actors assembling a tent outdoors. The participants exhibit varying levels of expertise and naturally make mistakes, which have been annotated. ATA [4] is a procedural dataset created for offline mistake detection in assembling activities. However, it provides only video-level mistake annotations, limiting its usefulness for frame-based applications. HoloAssist [3] provides egocentric videos of individuals performing multiple manipulation tasks following expert instructions. CaptainCook4D [28] is tailored for evaluating procedural activities, specifically focusing on the culinary domain. While comprehensive, [28] incorporates both procedural errors and attentional lapses. For our purposes, the latter category falls outside the scope of our investigation, yielding such dataset unsuitable. Notably,  [29] introduced a novel egocentric procedural error dataset consisting of videos depicting various errors within the cooking domain. In this study, we propose a new benchmark building upon PREGO tilize the datasets from [2, 1] as they provide insights into procedural errors in two distinct settings: controlled industrial and outdoor environments.
Table 1: Comparison of relevant models in procedural mistake detection. In the modalities column, RGB refers to RGB images, H stands for hand poses, E represents eye gaze, and K indicates keystep labels. Our approach is the first to adopt an egocentric, one-class and online method for detecting procedural mistakes. (*) denotes that these models are based on our protocol defined in [22].
@@
2.3 Methods for procedural mistake detection
Table 1 presents an overview of the core characteristics of recent methods for Mistake Detection in procedural videos. Ding et al. [5] propose a method for procedural learning that utilizes knowledge graphs to identify errors. Their approach is based on textual transcripts, ignoring the use of visual cues coming from videos.
In [4], the authors train an action recognition model and treat error detection as a semantic evaluation of the segmentation results. On the other hand, Assembly101 [2] and HoloAssist [3] apply the same error detection baselines on varying granularity but also run offline, requiring video segmentation and labels during training, thereby limiting their ability to detect previously unseen errors. In  [8], the authors address the problem of Procedure Step Recognition, which focuses on recognizing the correct completion of steps rather than the partial execution of activities. EgoPER [29] proposes a framework composed of an action segmentation module and a constrastive module for detecting errors unseen during training. PREGO [22] leverages video frames to detect procedural steps and incorporates symbolic reasoning for evaluating the procedure’s correctness. Its objective is to identify mistakes online as soon as they occur, ensuring prompt detection. Additionally, it is designed to function within an OCC framework, which allows it to detect a wide range of errors since it is trained on sequences without mistakes, without any constraint on specific errors  [20, 21].
Seminara et al. [16] learn task graphs by calculating the probability of witnessing a sequence of key steps. In contrast to the proposed methodology, which depends solely on context without further training for the anticipator, they employ a training approach grounded in Maximum Likelihood estimation.
@@ 2.4 Step recognition
Step recognition, the process of identifying discrete actions within a structured procedural sequence, is fundamental to progress in domains such as autonomous robotics and educational technology.
Zhuang et al. [13] introduce an action segmentation model that employs an attention-based architecture coupled with a Pairwise Ordering Consistency (POC) loss function. This approach effectively captures the correct sequence of procedural steps. Their innovation lies in developing a weakly-supervised method that requires only the set of actions in a procedure as input, eliminating the need for labor-intensive frame-level annotations. In Shah et al.[31], a novel loss for self-supervised learning is combined with a clustering algorithm to detect key steps in procedural videos without using any labels. [15] tackles the task by utilizing online instructional videos to learn actions and sequences without manual annotations, combining step recognition with a deep probabilistic model to account for step order and timing variability. An et al. [32] introduced MiniROAD, the state-of-the-art online step recognizer, which utilizes an RNN architecture and adjusts loss importance during training to perform online action recognition. However, existing methods like MiniROAD perform detection on a frame-by-frame basis, which can be limiting in practical scenarios that require recognizing sequences of actions rather than isolated frames. This limitation is particularly apparent in real-world situations where understanding the broader context of sequential actions is crucial. To overcome this, we introduce aggregation techniques to improve the coherence of action recognition, addressing the frame-level limitation inherent in current methods.
Moreover, the approach proposed in Shen et al. [33] also highlights the importance of holistically understanding actions rather than treating each frame in isolation. Yet, while PREGO focuses on detecting procedural mistakes using an OCC-based approach with LLMs for action anticipation, [33] targets action segmentation and progress estimation with task graphs.
@@ 2.5 Step Anticipation
Conversely, step anticipation predicts upcoming actions in a sequence, which is important for real-time AI decision-making. [34] generates multiple possible natural language outcomes, leveraging pretraining on a text corpus to handle the variability in future actions. [35] employs a two-level hierarchical approach, integrating both high-level human intentions and low-level action sequences, improving action anticipation in the long term. The pipeline consists of two models: the first extracts intentions and classifies actions, and the second generates future actions, conditioning human intentions to narrow down the uncertainty set of future actions.
Furthermore, the framework presented in [36] addresses future activity anticipation in egocentric videos by employing a contrastive loss to emphasize novel information and a dynamic reweighting mechanism that prioritizes informative past content. This approach improves video representation, leading to more accurate predictions of future activities. While we adopt a similar dual-branch architecture as in PREGO  [22]. However, we omit symbolic reasoning, instead leveraging the actual semantic labels to harness the reasoning capabilities of LLMs through an Automatic Chain of Thought.

### v2

@@ Abstract
Identifying procedural errors online from egocentric videos is a critical yet challenging task across various domains, including manufacturing, healthcare and skill-based training. The nature of such mistakes is inherently open-set, as unforeseen or novel errors may occur, necessitating robust detection systems that do not rely on prior examples of failure. Currently, no existing technique can reliably detect open-set procedural mistakes in an online setting. We propose a dual-branch architecture to address this problem in an online fashion: the recognition branch takes input frames from egocentric video, predicts the current action and aggregates frame-level results into action tokens while the anticipation branch leverages the solid pattern-matching capabilities of Large Language Models (LLMs) to predict action tokens based on previously predicted ones. Mistakes are detected as mismatches between the currently recognized action and the action predicted by the anticipation module.
Extensive experiments on two novel procedural datasets demonstrate the challenges and opportunities of leveraging a dual-branch architecture for mistake detection, showcasing the effectiveness of our proposed approach.
@@ 1 Introduction
Detecting procedural errors in videos has recently gained significant attention as it holds the potential to enhance safety and efficiency in several fields. Humans engage in procedural activities everyday in different contexts, such us daily activities (e.g., cooking a meal), hobbies (e.g., fixing a bike), and professional settings (e.g., operating complex machineries). Monitoring procedural activities is crucial to guarantee the desired results and operate safely, especially in high stakes contexts such as industrial workflows. Wearable devices, looking at the scene from the user’s point of view, and endowed with the capability of providing feedback through augmented reality, can be used to assess procedural mistakes (e.g., forgetting to add water to the kettle before turning it on) and provide real-time feedback. This immediate feedback allows for timely corrections, fostering faster skill development, skill acquisition and safer learning environments in high-stakes areas like surgery and industrial workflow.
@@
Figure 1: Procedural Mistake Detection involves identifying errors within a procedural video. Each procedure is composed of different steps that should be executed with a certain order. The aim is to develop a method capable of analyzing a video and determining whether each frame contains a mistake, e.g., a step that is not executed in the correct order. Addressing this task is particularly useful when applied to wearable devices, as they allow for direct feedback to the person performing the task. In the Figure, TI-PREGO, depicted by a gear icon, takes as input the video sequence from time 00 to $t-1$, classifying the current frame $t$ as either correct (green) or a mistake (red). This process continues frame by frame until a mistake is detected.
@@
Advanced mistake detection models will become essential to ensure precision, safety, and efficiency in procedural applications. This demand has recently resulted in an increasing number of datasets [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12] and methodologies designed to advance procedure learning [13, 4, 14, 15, 1, 6, 16] and error detection models [3, 5, 8]. However, existing mistake detection approaches vary widely. Some methods emphasize action detection, identifying specific errors like missing steps, out-of-sequence actions [17] or incorrect ordering [16], while other models bypass action detection altogether and monitor changes to the assembled object to ensure procedural accuracy [18, 19]. This variability in proposed approaches and their intended objectives highlights the need for more comprehensive evaluation frameworks and standardized benchmarks to assess and compare the effectiveness of different mistake detection strategies and for more olistic models, able to detect any deviation from the correct workflow.
@@
An ideal Procedural Mistake Detection (PMD) model should demonstrate two critical capabilities: robustness to diverse mistake types and the ability to provide online, timely feedback. Regarding robustness, a reliable PMD system should be capable of identifying any mistake that may occur within a structured procedure, regardless of its type. To uphold this principle, in this paper we propose framing mistake detection as a One-Class Classification problem, where the model is trained to learn the correct sequences of steps within a procedure and is then able to recognize any deviation from the correct workflow. Within the OCC framework, during training, the PMD model is only exposed to “normal” procedures [20, 21] to learn the correct workflows, i.e. the sequences of performed steps that allow to correctly complete the task. Conversely, during inference, the learned model should detect any deviation from the learnt workflow, deeming that step as an error. Further, it is worth noting that the OCC framework requires only binary labels (Correct vs. Mistake), which are easier to collect with respect to fine-grained labels like the error type when labelling videos of procedural datasets. Second, the model should provide online [^1] We distinguish an online setup, which allows slight processing delays while enabling effective feedback, from a real-time setup, which requires immediate responses. feedback quickly enough to allow users to correct mistakes promptly, minimizing the reinforcement of incorrect actions and reducing the risk of hazardous situations. Besides these two pivotal abilities, the PMD system should leverage an egocentric perspective in procedural videos, enhancing the model’s applicability by mirroring natural human perception, and allowing for seamless real-world deployment on wearable devices.
@@
To conform to these requirements, we propose a dual-branch approach for detecting mistakes in procedural videos sequentially combining step recognition and anticipation and identifying mistakes as mismatches in predicted and observed actions. The first branch, the step recognition module, processes the input video online and classifies the current action being performed. The second branch, the step anticipation module, predicts the next step to perform based on the steps recognized by the step recognition module. We implement the anticipator branch as a Large Language Model (LLM), leveraging several prompting schemes to interrogate it and retrieve the next action to be performed. The model will then flag a step as mistaken when there is a divergence between recognized and anticipated actions, highlighting inconsistencies between the action executed by the operator and the expected steps.
@@
Our contributions include (1) a novel benchmark for online mistake detection, encompassing two dataset (Assembly101-O and Epic Tent-O), (2) exploring and benchmarking multiple LLMs for the task of step anticipation aiming to select the most robust candidate for our anticipator module, (3) experimenting with various prompting techniques, including Automatic Chain of Thought (Auto-CoT), which enhances overall performance, reporting state-of-the-art results on Assemblty101-O and Epic Tent-O, (4) investigating different prediction aggregation techniques to improve mistake detection robustness and continuity. Extensive experimentation demonstrates the challenges and opportunities in dual-branch architectures and LLM-based step anticipation for open-set mistake detection.
This paper extends our previous conference work, PREGO [22], by refining key components and methods to improve procedural mistake detection (see Fig. 1). Unlike our original study, which primarily reported classical Precision and Recall metrics, in this work, we adopt Balanced Precision [23] to mitigate the extreme class imbalance inherent in egocentric mistake detection. Balanced Precision scales the False Positives by the ratio of positive‐to‐negative samples, ensuring that the penalty for mislabeling a correct step is placed on the same scale as the reward for correctly identifying a mistake. Specifically, we lead a comprehensive evaluation of several LLMs as step anticipators, increasing the state‑of‑the‑art results on Assembly101‑O and Epic Tent‑O [22]. We compare a broad range of LLMs such as Llama‑2 [24], Llama‑3 [25], Gemma [26], and Mistral [27], fine‑tune them using Low‑Rank Adaptation (LoRA), and experiment with prompting methods including Zero‑Shot, Few‑Shot, and Auto‑CoT [28].
Further, we revisit the recognizer branch’s frame-by-frame prediction mechanism, aiming to mitigate inconsistencies and reduce noise in its predictions. To deal with this, we extend beyond fixed-length window aggregation to explore alternative strategies that reduce noise and improve prediction accuracy.
@@ 2 Related Works
This section reviews the foundations of current approaches related to online detection of procedural mistakes,. We first survey existing models and compare them across key design dimensions (Sec. 2.1). We then analyze the most relevant procedural video datasets (Sec. 2.2), examine step recognition (Sec. 2.3) and anticipation (Sec. 2.4) methods. Finally, we discuss how recent advances in large language models enable symbolic reasoning (Sec. 2.5). Together, these components define the landscape for building effective online, open-set mistake detectors.
@@ 2.1 Procedural Mistake Detection
Procedural learning has seen notable progress due to the development of diverse datasets [6, 10, 9, 11, 12], which offer valuable insights into both structured [7, 3, 29] and unstructured tasks [1, 30]. These datasets span a wide range of applications—from industrial assembly [29, 2, 8, 12] to everyday cooking activities [30, 31, 32]—yet the lack of a standardized benchmark for mistake detection has led to limited evaluation and a fragmented literature.
In Table 1, we compare relevant models for procedural mistake detection across multiple dimensions: whether they operate in an open-set or one-class condition, whether they support online or offline detection, the data modalities they leverage (e.g., RGB frames, hand poses, eye gaze, keystep logic), the task they address, and the datasets used. Notably, models marked with an asterisk (*) follow our protocol introduced in [22], standardizing an online, open-set detection scenario for consistent evaluation.
Existing methods illustrate a variety of approaches. For instance, Ding et al. [5] employ knowledge graphs with textual transcripts, neglecting visual cues. Ghoddoosian et al. [4] focus on action recognition, treating error detection as a semantic evaluation of segmentation results. Other works, such as Assembly101 [2] and HoloAssist [3], apply error detection baselines offline, relying on pre-segmented videos and annotated labels. By contrast, [8] emphasizes Procedure Step Recognition, focusing on correct step completion rather than partial activities, while EgoPER [33] integrates action segmentation with contrastive learning to detect previously unseen errors. Our prior work, PREGO [22], uses RGB frames alongside symbolic reasoning to assess procedural correctness online, operating under an OCC framework and trained solely on error-free sequences [20, 21]. Finally, Seminara et al. [16] adopt a distinct approach, learning task graphs via Maximum Likelihood estimation of key-step sequences.
In contrast to prior approaches that rely on offline processing, pre-segmented videos, or textual-only cues, our method uniquely fuses real-time visual analysis with symbolic reasoning within an OCC framework. By learning solely from error-free sequences, our approach can detect known and unforeseen mistakes without being constrained to predefined error types. As indicated in Table 1, we are the first to propose a fully online OCC protocol , allowing detection to be performed frame-by-frame without requiring prior action segmentation. This design enables a more robust, context-aware, and prompt detection of errors, addressing the key limitations observed in existing literature.
@@ 2.2 Datasets of Procedural Egocentric Videos
Previous investigations proposed datasets and benchmarks to support research in procedural mistake detection. IndustReal [8] focuses on a single toy, resulting in the learning of only one procedure without annotations for procedural mistakes. Assembly101 [2] is a large-scale video dataset that provides frame-level mistake annotations. It features videos of actors assembling toys, with synchronized Ego-Exo views and annotated hand positions. IndustReal [8] focuses on a single toy, resulting in the learning of only one procedure without annotations for procedural mistakes. Epic-tent [1] is a dataset that covers a different domain of unscripted actions that capture actors assembling a tent outdoors. The participants exhibit varying levels of expertise and naturally make mistakes, which have been annotated. ATA [4] is a procedural dataset created for offline mistake detection in assembling activities. However, it provides only video-level mistake annotations, limiting its usefulness for frame-based applications. HoloAssist [3] provides egocentric videos of individuals performing multiple manipulation tasks following expert instructions. CaptainCook4D [34] is tailored for evaluating procedural activities, specifically focusing on the culinary domain.
Notably,  [33] introduced a novel egocentric procedural error dataset consisting of videos depicting various errors within the cooking domain.
In this study, we propose a new benchmark building upon the datasets from [2, 1] as they provide insights into procedural errors in two distinct settings: a controlled industrial environment and outdoor environments.
Table 1: Comparison of relevant models in procedural mistake detection. In the modalities column, RGB refers to RGB images, H stands for hand poses, E represents eye gaze and K indicates keystep labels. Our approach is the first to adopt an egocentric, one-class and online method for detecting procedural mistakes. (*) denotes that these models are based on our protocol defined in [22].
@@ 2.3 Step Recognition
Step recognition is the process of identifying discrete actions within a structured procedural sequence.
Zhuang et al. [13] introduce an action segmentation model that employs an attention-based architecture coupled with a Pairwise Ordering Consistency (POC) loss function. Moreover, they develop a weakly-supervised method that requires only the set of actions in a procedure as input, eliminating the need for labor-intensive frame-level annotations. In Shah et al. [36], a novel loss for self-supervised learning is combined with a clustering algorithm to detect key steps in procedural videos without using any labels. [15] tackles the task by utilizing online instructional videos to learn actions and sequences without manual annotations, combining step recognition with a deep probabilistic model to account for step order and timing variability.
An et al. [37] introduced MiniROAD, the state-of-the-art online step recognizer, which leverages an RNN architecture and adjusts loss importance during training to perform online action recognition. The approach proposed in Shen et al. [38] also highlights the importance of holistically understanding actions rather than treating each frame in isolation. In contrast to previous methods that rely on frame-by-frame detection—such as MiniROAD [37], which focuses on step recognition—our approach addresses the limitations of isolated frame-level analysis. While prior work has established a strong foundation in procedural step recognition, it often falls short in capturing the broader temporal context that is critical in real-world scenarios. To overcome this, we propose and evaluate several aggregation techniques that enhance the coherence of action recognition across sequences, thereby enabling more robust and context-aware mistake detection.
@@ 2.4 Step Anticipation
Step anticipation is the process of predicting the next action in a procedural sequence before it occurs.
[39] generates multiple possible natural language outcomes, leveraging pretraining on a text corpus to deal with the variability in future actions. [40] employs a two-level hierarchical approach, integrating both high-level human intentions and low-level action sequences, improving action anticipation in the long term. The pipeline consists of two models: the first one extracts intentions and classifies actions while the second one generates future actions, conditioning human intentions to narrow down the uncertainty set of future actions. The framework presented in [41] addresses future activity anticipation in egocentric videos by employing a contrastive loss to emphasize novel information and a dynamic reweighting mechanism that prioritizes informative past content. This approach improves video representation, leading to more accurate predictions of future activities.
Similar to AntGPT [42], we leverage video frames to anticipate future actions using an LLM enhanced with Automatic Chain of Thought. However, while previous approaches, such as AntGPT, require predefined segmentation of a long video into ordered, annotated short segments with corresponding action labels, our online setting processes each frame independently without prior segmentation.